OpenStudy (anonymous):

The number of gears G a machine can make varies directly as the time T of operation. If it can make 610 gears in 2 hours, how many gears can it make in 5 hours?

OpenStudy (anonymous):

5/2 = 2.5 2.5 x(times) 610 = 1220 + 305 = 1525

OpenStudy (lilg132):

610 / 2 = Will give you how many gears it can do in 1 hour so 610 / 2 = 305 so now we know it makes 305 gears in 1 hour now to work out how many gears it makes in 5 hours we just multiply it by 5 so 305 * 5 = 1525 gears are made in 5 hours dont forget to click good answer :)